OpenStudy (kayla2001):

Joanna went school supply shopping. She spent $35.72 on notebooks and pencils. Notebooks cost $2.24 each and pencils cost $1.48 each. She bought a total of 19 notebooks and pencils. How many of each did she buy?

OpenStudy (kayla2001):

hey there every question you answer ill give you a medal deal?

OpenStudy (anonymous):

fan too?

OpenStudy (kayla2001):

im already yoour fan

OpenStudy (anonymous):

oh haha 1 sec

OpenStudy (kayla2001):

ok

OpenStudy (sleepyjess):

We need to make 2 equations to represent this. Do you know how to do that?

OpenStudy (anonymous):

Pretty sure its 11 and 8

OpenStudy (anonymous):

11 notebooks

OpenStudy (anonymous):

medal and @ for more

OpenStudy (texaschic101):

n + p = 19 ----> n = 19 - p 2.24n + 1.48p = 35.72 now sub 19 - p in for n in the 2nd equation 2.24(19 - p) + 1.48p = 35.72 42.56 - 2.24p + 1.48p = 35.72 -2.24p + 1.48p = 35.72 - 42.56 - 0.76p = - 6.84 p = -6.84/-0.76 p = 9 n + p = 19 n + 9 = 19 n = 19 - 9 n = 10 check.. 2.24n + 1.48p = 35.72 2.24(10) + 1.48(9) = 35.72 22.40 + 13.32 = 35.72 35.72 = 35.72 (correct) notebooks (n) bought were 10 and pencils (p) bought were 9